The difference of two integers is 46.If the larger integer is twice the smaller.Find the integers
Answer by jsmallt9(3759)

Let x = one integer and let y = the other integer. Then
x - y = 46
and
x = 2y

There are several ways to solve systems of equation like this. Since the second equation is already solved for x, the Substitution Method looks easiest. So we'll substitute the expression for x, 2y, in the second equation for the x in the first equation:
(2y) - y = 46.
(Note the use of parentheses. It is an extremely good habit to use parentheses like this whenever you substitute one expression for another. In this problem it doesn't make much difference but in many problems it is critical.)
Subtracting we get:
y = 46
Now that we have y we can find x. We can use x = 2y with y = 46:
x = 2(46) = 92

So the two integers are 46 and 92.